### v1

# EntGPT: Linking Generative Large Language Models with Knowledge Bases

Abstract.
The ability of Large Language Models (LLMs) to generate factually correct output remains relatively unexplored due to the lack of fact-checking and knowledge grounding during training and inference. In this work, we aim to address this challenge through the Entity Disambiguation (ED) task. We first consider prompt engineering, and design a three-step hard-prompting method to probe LLMs’ ED performance without supervised fine-tuning (SFT). Overall, the prompting method improves the micro-$F_{1}$ score of the original vanilla models by a large margin, on some cases up to 36% and higher, and obtains comparable performance across 10 datasets when compared to existing methods with SFT. We further improve the knowledge grounding ability through instruction tuning (IT) with similar prompts and responses. The instruction-tuned model not only achieves higher micro-F1 score performance as compared to several baseline methods on supervised entity disambiguation tasks with an average micro-$F_{1}$ improvement of 2.1% over the existing baseline models, but also obtains higher accuracy on six Question Answering (QA) tasks in the zero-shot setting. Our methodologies apply to both open- and closed-source LLMs.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Question answeringccs: Information systems Information extractionccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generation
1. Introduction
Figure 1. Pipeline of the proposed frameworks: Given input document with the annotated mention, both the approaches first conducts (1) entity candidate generation step to obtain relevant entities. EntGPT-Prompting (EntGPT-P) then performs (2) an augmentation step to obtain an auxiliary content of the annotated mention followed by (3) a multi-choice selection prompt to decide the corresponding entity of the annotated mention. EntGPT-Instruction tuning (EntGPT-I), on the other hand, goes through the (2) construction of the instruction dataset, and is (3) instruction-tuned on the curated data.
The introduction of Large Language Models (LLMs) has sparked intense interest and scrutiny regarding tools that leverage their capabilities. Although their text-generation and unstructured reasoning capabilities are “remarkable” (Raffel et al., 2020; Radford et al., 2018, 2019; Brown et al., 2020; Touvron et al., 2023b), their ability to produce structured output remains underdeveloped and relatively unexplored (Yu et al., 2022; Sun et al., 2023). Moreover, tools leveraging the LLMs are known to generate content that is factually incorrect and nonsensical (Ji et al., 2023). This phenomena, commonly known as hallucination, occurs, at least partly, due to a lack of factual knowledge grounding during pre-training. To address this issue, researchers have introduced various strategies such as encompassing feedback mechanisms, external information retrieval, and early refinement in language model generation (Tonmoy et al., 2024). However, these models typically do not ground the generated knowledge to well-structured knowledge bases. In the present work, we mitigate these common problems by linking (i.e., grounding) keyterms to their entries in a knowledge base and using that structured knowledge to support the text generation task.
The entity disambiguation (ED) task is a type of information extraction (IE) that seeks to link text fragments that represent some real-world entity with an entry in some structured knowledge base like an encyclopedia or a dictionary. Once linked, the existing data and its relationships within the knowledge base could be used to assist in a number of downstream processes. The integration of deep neural network models, such as Transformers (Vaswani et al., 2017), utilized in LLMs, with the symbolic reasoning capabilities of existing knowledge bases, has been dubbed as the Third Wave of AI (DARPA, 2018) by the United States Defense Advanced Research Projects Agency (DARPA). This convergence is considered essential for the advancement of problem solving and reasoning in AI systems (Brooks, 1981; Hitzler and Sarker, 2022). For such a vision to become reality, it is critical that LLMs be linked to the symbolic entities that they reference.
In the present work and as illustrated in Fig. 1, we explore two distinct approaches to ground the generative language model through the ED task. The first is a multi-step prompting framework, which emulates prompt-tuning called Entity GPT-Prompting (EntGPT-P). This approach first generates a set of entity candidates for the mentions in the document. Then, it utilizes the power of the LLM to generate auxiliary content facilitating the final selection of the corresponding entity from the candidate set. Results over various tasks and datasets show that EntGPT-P achieves comparable performance to supervised models even without training/fine-tuning on human-annotated data.
Secondly, we implemented an Instruction Tuning (IT) approach , named EntGPT-Instruction tuning (EntGPT-I), which is further trained on instructions (i.e., Prompt-Response pairs) and exhibits more controllable and predictable behavior (Zhang et al., 2023b). EntGPT-I obtained significantly better results (micro-$F_{1}$ score) compared to EntGPT-P across ED benchmarks. We also evaluated EntGPT-I on six closed-world question answering tasks. Experiments showed that EntGPT-I achieved state of the art (SOTA) performance (micro-$F_{1}$ score) in a zero-shot setting while boosting accuracy scores over the most popular and powerful commercial LLMs.
Our major contributions can be summarized as follows:
  • We introduce EntGPT-P, a model that follows a three-step hard prompting method to guide generative language models. Our findings demonstrate that EntGPT-P achieves comparable, and sometimes better performance when compared to supervised counterparts across different benchmarks on the ED task.
  • We present a model instruction-tuned on generative LLMs(i.e., EntGPT-I) with (instruction - response) pairs for the ED task, while augmenting the instructions with enhanced knowledge. We achieve SOTA performance on the ED task.
  • We show that these knowledge-enhanced generative models help to generate factually correct answers compared to commercial LLMs through six different Question Answering tasks. EntGPT-I achieves SOTA performance on all datasets in zero-shot setting and comparable performance to supervised methods.
  • We also provide a detailed performance comparison of our proposed approaches on both open-source (e.g., Llama2) and closed-source (e.g., GPT-3.5) LLMs. We find a consistent boost in micro-$F_{1}$/accuracy performance when compared to their vanilla counterparts.
Throughout the remainder of the paper, unless explicitly stated otherwise, we use the following abbreviations: GPT3.5 for gpt-3.5-turbo, Llama for Llama-65B, Llama2 for Llama2-70B, and Llama2-chat for its chat variant Llama2-70b-chat.
2. Related Work
2.1. Hallucination in LLMs
The so-called hallucination problem has impeded the use of the intelligent systems that leverage LLMs in sensitive applications such as summarizing financial analysis reports, medical records, and customer support interactions (Tonmoy et al., 2024). To overcome the challenges raised by the generation of inaccurate and erroneous information, researchers have proposed several techniques. In a recent survey paper, Tonmoy et. al (Tonmoy et al., 2024) categorized these techniques into two different subgroups: Prompt Engineering, and Models Development. Prompt engineering explores generating more factual and grounded output with well-curated instructions/context to the model; model development requires algorithmic and/or architectural improvement.
Within the prompt engineering paradigm, researchers have explored the use of external knowledge to provide updated and more-thorough context to the model (Peng et al., 2023; Varshney et al., 2023; Lewis et al., 2020), refinements to the generated output through proper feedback to the model (Si et al., 2022; Dhuliawala et al., 2023; Lei et al., 2023; Mündler et al., 2023), and prompt tuning (Cheng et al., 2023; Jones et al., 2023). Moreover, techniques involving decoding strategies (Shi et al., 2023; Chuang et al., 2023; Li et al., 2023), use of of Knowledge Graphs (KGs) (Ji et al., 2022; Bayat et al., 2023), faithfulness-based loss functions (Yoon et al., 2022; Qiu et al., 2023), and supervised fine-tuning (Elaraby et al., 2023; Tian et al., 2023; Zhang et al., 2023a) have been introduced to tackle hallucination through the algorithmic approach. The EntGPT-P method described in the present work falls under the prompting engineering paradigm, whereas the EntGPT-I model tackles hallucination through an instruction-tuning approach.
2.2. Entity Disambiguation
Existing research in ED is primarily divided into two categories: feature extraction and task refinement. Key work in feature extraction includes the interaction between mentions and context (Ganea and Hofmann, 2017; Kolitsas et al., 2018), the consistency between entities and entity types (Raiman and Raiman, 2018; Tedeschi et al., 2021; Raiman, 2022), the relationship between entities and knowledge base entries (Ayoola et al., 2022a, b), and correlation between entities (Phan et al., 2019; Xiao et al., 2023b). Traditional ED models framed this task as entity classification (Ganea and Hofmann, 2017), while more recent work approached it as machine reading comprehension (Barba et al., 2022a), dense retrieval (Wu et al., 2020), question answering (Zhang et al., 2022b), sequence-to-sequence generation (Cao et al., 2021), and prompt LLMs (Xiao et al., 2023a).
2.3. Prompt-based Methods for Information Extraction
One of the important developments in the field of prompt-based methods for information extraction was the release of GPT-3, which has been shown to outperform state-of-the-art methods for language modeling (Brown et al., 2020). The benefit of prompt-based methods is the ability to query a language model to obtain additional knowledge and understanding about a given context, thereby significantly improving performance in zero-shot and few-shot settings (Wang et al., 2023b; Sun et al., 2023).
There are mainly three streams of work in this paradigm. The first is in-context learning where a few demonstrative examples are provided to an LLM in order to obtain some desired output. The second is prompt-tuning, where a set of learnable parameters are inserted into the model (Lester et al., 2021; Ding et al., 2022) or each layer (Li and Liang, 2021; Wang et al., 2023a) for fine-tuning. The third line of work is instruction-tuning, where (instruction, response) pairs are prepared used to fine-tune the LLM (Longpre et al., 2023).
In terms of task formulation, existing work on prompt-based methods for information extraction can be primarily categorized into three paradigms: type-oriented, span-oriented, and generation. Type-oriented tasks (Ding et al., 2022) mainly locate the given class for a certain mention within the original documents. Span-oriented methods (Cui et al., 2021) enumerate the possible spans and assign corresponding class labels. Finally, generation aims to transform information extraction into a sequence generation schema (Wang et al., 2023b).
3. Methodology
Here, we formally define the ED task and describe the general pipeline employed in both the prompt-based and instruction-tuning approaches to ground generative LLMs with entities from a knowledge base.
3.1. Problem Definition
Given a document represented as a sequence of tokens $\mathcal{D}$ and a set of subsequences $M=\{m_{1},\cdots,m_{n}\}$ (i.e., mentions) contained in $\mathcal{D}$. The goal of the ED task is to establish a mapping for each mention $m_{i}\in M$ to its corresponding entity $e$ in the set $\mathcal{E}$ representing entities in some knowledge base (such as Wikipedia or DBpedia).
3.2. EntGPT-P: Entity Disambiguation with Multi-Step Prompting
We view the task of grounding (or linking) generative LLMs to knowledge bases as a natural use case of ED where text sequences need to be matched with the relevant knowledge base entries to permit the knowledge base’s well-curated structure to be leveraged for downstream tasks.
The EntGPT-Prompting (EntGPT-P) model introduced in this section formulates the ED task as having three steps: 1) Generate and filter entity candidates from knowledge base $\mathcal{E}$; 2) Augment each mention by extracting relevant information with a prompt for a LLM; 3) Combine candidates from Step 1 with context from Step 2, forming a multi-choice question prompt for a LLM.
3.2.1. Step 1: Entity Candidate Generation
Given a mention $m$ in document $D$, the first step is to find a subset of entity candidates $\mathcal{E}_{c}$ corresponding to $m$ within the knowledge base $\mathcal{E}$. We combine two strategies to generate entity candidates. First, we use statistical information of hyperlinks in the knowledge base to surface the most likely entity candidates $\mathcal{E}_{p}$ from $\mathcal{E}$ that exhibit syntactic similarity to the mentions within the document (Ganea and Hofmann, 2017). However, this probabilistic approach, we call the Prior probability-approach or simply Prior, suffers from low recall because it only considers syntactic similarity. In low-recall cases, we additionally employ a dense retrieval model as a backup to augment the entity candidate generation. Specifically, we use the Bidirectional Encoder Representations from Transformers (BERT)-based entity linking (BLINK) model, which primarily uses cleaned Wikipedia hyperlinks to as its retrieval model (Wu et al., 2020), to generate extra entity candidates $\mathcal{E}_{r}$. Therefore, the final entity candidates set from step 1 is $\mathcal{E}_{c}=\mathcal{E}_{p}\cup\mathcal{E}_{r}$ from which we select the top-10 candidates.
3.2.2. Step 2: Augmentation by Prompting
Because many candidates are syntactically similar, distinguishing them is challenging without contextual information. So we augment mentions with relevant information by asking the LLM “What does {mention} represent?” For example, from Fig. 1 (at left), we generate auxiliary content $\mathcal{A}$ of George Bush by asking the LLM “What does George Bush represent?”. We found that generating context information in this way has the following two advantages: 1) The content generated by the LLM considers contextual information from $\mathcal{D}$ and is supplemented with the world knowledge encoded within the LLM. 2) The auxiliary content produced by the system is more specifically targeted towards our task and therefore minimizes the impact of superfluous information.
3.2.3. Step 3: Multiple-choice Selection by Prompting
Armed with the entity candidates set $\mathcal{E}_{c}$ (from Step 1) and the auxiliary content $\mathcal{A}$ (from Step 2), the goal of Step 3 is to select the most appropriate $e\in\mathcal{E}_{c}$. As illustrated in Fig. 1 (at left), we employ an instruct-based prompt to direct the LLM to make a selection from $\mathcal{E}_{c}$ using the auxiliary content $\mathcal{A}$. To distinguish among these entity candidates effectively, we include the first sentence extracted from the knowledge base entry of each entity candidate as a descriptive reference. Also note that each of the entity candidates come from the subset $\mathcal{E}_{c}$ obtained in Step 1, rather than the complete knowledge base $E$. Therefore, this multi-choice setting can not address the situation where the correct entity is not included in $\mathcal{E}_{c}$. To accommodate this situation, we include the option “None” as a catch-all choice for handling such cases.
3.3. EntGPT-I: Entity Disambiguation with Instruction Tuning
In an initial study, described below, we found that EntGPT-P substantially reduces hallucination and helped to deliver more accurate results compared to non-augmented foundation models (see metrics in Tables 2 and  3). Nonetheless, EntGPT-P still under-performed supervised methods indicating room for improvement. To that end, we constructed an additional instruction-tuning model on top of generative LLMs through similar prompts as that of EntGPT-P. In this section, we discuss the processes involved in creating the EntGPT-I model including the construction of an instruction dataset and the instruction-tuning procedure.
3.3.1. Instruction Dataset Construction
An instruction dataset usually consists of three different elements: an instruction, an optional input, and an anticipated output (Ji et al., 2023). Usually in the form of natural language, an instruction provides the description of the task to the model, whereas any optional input can provide additional context. In the present work, we curate one of the largest manually annotated entity linking and entity disambiguation datasets, i.e., Accurate online Disambiguation of Entities in Text and Tables (AIDA) (Hoffart et al., 2011) into the required format (instruction, input, output) needed to fine-tune the an LLM model. An example of such a training instance is shown in Fig1 (at right).
3.3.2. Instruction Tuning
We fine-tune two generative models (GPT3.5 and Llama2-chat) in a fully supervised manner on the curated instruction dataset. We used the OpenAI API platform (Kublik and Saboo, 2023) to fine-tune GPT3.5, and we used Replicate API [^1] https://replicate.com/docs/get-started/python to fine-tune Llama2-chat. Fine-tuning was performed for only one epoch and we kept all other hyperparameters at their default settings. Although hyper-parameter tuning could potentially enhance performance, hyper-parameter search is outside the scope of this study. After fine-tuning, EntGPT-I is then evaluated on both the ED task and the question answering task.
4. Experiments
To test the two models described in the present work, we used ten public entity disambiguation benchmark datasets and six question answering datasets widely used in the literature. The entity disambiguation task was to link mentions in the dataset with the correct entity in a knowledge base; the question answering task was to determine the correct answer from multiple choices. Using these experiments our goal is to address the following research questions (RQs):
  • RQ1: Does multi-step prompting (EntGPT-P) and instruction tuning (EntGPT-I) produce more accurate output (i.e., reduce hallucination)?
  • RQ2: Does fine-tuning generative LLMs with auxiliary information gleaned from the linked knowledge base improve accuracy?
  • RQ3: How do the various steps involved in the model contribute to the performance of EntGPT?
  • RQ4: Do correlations among entities in the knowledge base help to mitigate hallucinations in large language models?
To answer these questions directly, we first introduce the datasets that we will use and several SOTA models.
4.1. Datasets
Following general ED settings introduced in prior work (Guo and Barbosa, 2018; Kolitsas et al., 2018; Ayoola et al., 2022b; Barba et al., 2022a), we evaluated our models on six ED datsets: AIDA (Milne and Witten, 2008), MSNBC (MSN) (Cucerzan, 2007), Aquaint text corpus (AQUAINT/AQU) (Milne and Witten, 2008), a multilingual training corpus ACE2004 (ACE04) (Ratinov et al., 2011), Walking Named Entity Disambiguation (WNED-WIKI) (Guo and Barbosa, 2018), and WNED-CWEB (CWEB) (Gabrilovich et al., 2013). We also consider five additional datasets from GERBIL (Usbeck et al., 2015): Keyphrase Overlap Relatedness for Entity Disambiguation (KORE50/KORE) (Hoffart et al., 2012), Open knowledge extraction challenge (OKE-2015/OKE15) (Nuzzolese et al., 2015), OKE-2016 (OKE16) (Nuzzolese et al., 2016), N3-Reuters-128 (REU) (Röder et al., 2014), and N3-RSS-500 (RSS) (Röder et al., 2014). We provide the statistics of these benchmarks, including the number of documents, and mentions in Table 1. Each of the experiments used Wikipedia as the background knowledge base. Following the example of Guo and Barbosa (2018), we manually removed spurious mentions that do not appear in the knowledge base as well as repeated documents and empty documents without any mentions.
We found that EntGPT-I (with GPT3.5 as the foundation model) outperformed other approaches on the average micro-F1 score (Tables 2 and  3), so we additionally evaluated EntGPT-I (GPT3.5) on six different commonsense reasoning (question answering) benchmarks: AI2 Reasoning Challenge-Challenge Set (ARC-C) (Clark et al., 2018), OpenbbookQA (OBQA) (Mihaylov et al., 2018), Abductive Natural Language Inference (aNLI) (Bhagavatula et al., 2019), CommonsenseQA (CSQA) (Talmor et al., 2018), Physical Interaction QA (PIQA) (Bisk et al., 2020), and Social Interaction QA (SIQA) (Sap et al., 2019).
4.2. Comparison Models
For ED, we categorize the comparison models into three different groups according to their fine-tuning data: Wikipedia and AIDA (Wiki+AIDA), AIDA-only (AIDA), and no fine-tuning (zero-shot). We compared the performance of our models against the following models:
  • End2End (Kolitsas et al., 2018): The End-to-End Neural Entity Linking (End2End) is a global ED and entity linking system based on Deep-ED (Ganea and Hofmann, 2017), but with a more robust recurrent neural network (RNN) architecture that considers mention-context and mention-mention interactions.
  • REL (van Hulst et al., 2020): The Radboud Entity Linker (REL) is an ED and entity linking package which combines the NER method FLAIR (Akbik et al., 2019) and ED method mulrel-nel (Le and Titov, 2018).
  • GENRE (Cao et al., 2021): The Generative Entity REtrieval model (GENRE) is a generation-based model that re-casts the ED task as an entity name generation process.
  • ner4el (Tedeschi et al., 2021): The Named Entity Recognition for Entity Linking (NER4EL) model uses named entity recognition to perform the entity linking task, which is similar to the entity disambiguation task.
  • ReFinED (Ayoola et al., 2022b): The Representation and Fine-grained typing for Entity Disambiguation (ReFinED) model is an efficient end-to-end zero-shot-capable EL model is a recent system that considers Wikipedia entries as extra features to the ED and entity linking tasks.
  • ExtEnd (Barba et al., 2022b): The Extractive Entity Disambiguation (ExtEnd) model is a transformer-based framework where the disambiguation task is recast as a machine reading comprehension task.
  • CoherentED (Xiao et al., 2023b): The Coherent Entity Disambiguation (CoherentED) model aims to enhance the coherence of entity predictions and achieves state-of-the-art performance on most datasets.
In the question answering task we categorize the baselines into two different groups: supervised and zero-shot. In the supervised setting, we follow common experimental methods to report the performance for the following models.
  • ROBERTa (Liu et al., 2019): The Robustly Optimized BERT Pretraining Approach (ROBERTa) is an optimized version of the Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2018) model, which uses Masked Language Modeling (MLM) as its pre-training objective. The pre-trained model is fine-tuned on each question answering benchmark in a fully supervised manner.
  • QAGNN & GreaseLM (Yasunaga et al., 2021; Zhang et al., 2022a): The Question Answering Graph Neural Network (QAGNN) and the Graph REASoning Enhanced Language Model (GreaseLM) both fine-tune an LLM using a knowledge graph (KG) on downstream tasks. These architectures are broadly referred to as KG-augmented question answering models.
  • DRAGON (Yasunaga et al., 2022): Deep Bidirectional Language-Knowledge Graph Pretraining (DRAGON) uses a knowledge base to enhance the pre-training stage and also performs fine-tuning on downstream tasks in a supervised manner.
In the zero-shot setting, we use the following LLMs as foundation models: Large Language Model Meta AI (Llama, Llama2, & Llama2-chat(Touvron et al., 2023a, b), the Fine-tuned LAnguage Net (FLAN-137B) (Wei et al., 2021), and the Generative Pre-trained Transformer (GPT3.5) (Ouyang et al., 2022).
Table 1. Data statistics for 11 experimental datasets. Number of documents (# of Docs), number of mentions (# of Mention).
AIDA MSN AQU ACE04 CWEB WIKI KORE OKE15 OKE16 REU RSS
# of Docs 231 20 50 35 320 319 50 101 173 113 357
# of Mention 4485 656 727 257 11154 6793 144 536 288 650 524
4.3. Evaluation Metrics
To maintain a fair comparison across datasets, we use the in-KB micro-F1 score as our evaluation metric for ED following the example of Guo and Barbosa (2018). Specifically, being in-KB requires that ground truth mentions correspond to existing knowledge base entries. Empty or out-of-KB entities are removed in the evaluation process. The mean Micro-F1 score is taken per-mention. Although a model may predict non-entities, each mention will have some corresponding ground truth entity. For question answering, we use accuracy as the evaluation metric.
4.4. Results and Findings
In this section, we discuss the results and findings from our study on both the entity disambiguation and question answering benchmarks.
4.4.1. Entity Disambiguation (RQ1 and RQ2)
Table 2. Test micro-F1 scores on standard six benchmarks. With each setting, best scores are highlighted in bold. † indicates that the results are obtained directly from the source paper.
Fine-tuning Data Model AIDA MSN AQU ACE04 CWEB WIKI avg
Wiki+AIDA REL (van Hulst et al., 2020) 0.928 0.935 0.873 0.897 0.776 0.780 0.864
End2End (Kolitsas et al., 2018) 0.891 0.933 0.894 0.892 0.760 0.740 0.851
GENRE (Cao et al., 2021) 0.933 0.943 0.899 0.901 0.773 0.874 0.887
ReFinED (Ayoola et al., 2022b) 0.939 0.941 0.918 0.908 0.794 0.874 0.894
ExtEnD (Barba et al., 2022a) 0.926 0.947 0.916 0.918 0.777 0.888 0.895
CoherentED (Xiao et al., 2023b) 0.894 0.963 0.946 0.934 0.811 0.906 0.909
AIDA GENRE (Cao et al., 2021) 0.886 0.881 0.771 0.823 0.719 0.717 0.795
ner4el (Tedeschi et al., 2021) 0.925 0.892 0.695 0.913 0.685 0.640 0.792
ExtEnd (Barba et al., 2022a) 0.900 0.945 0.879 0.889 0.766 0.767 0.858
EntGPT-I (Llama2) 0.691 0.744 0.583 0.681 0.556 0.604 0.643
EntGPT-I (GPT3.5) 0.920 0.922 0.906 0.937 0.795 0.791 0.879
Zero-Shot Llama2 0.432 0.401 0.441 0.381 0.347 0.451 0.408
EntGPT-P (Llama2) 0.708 0.741 0.635 0.746 0.584 0.632 0.674
GPT3.5 0.746 0.821 0.753 0.853 0.638 0.684 0.749
EntGPT-P (GPT3.5) 0.821 0.867 0.791 0.918 0.709 0.808 0.819
A performance comparison over six entity disambiguation benchmarks is presented in Table 2. We categorized these experiments into three groups based on the data used during the fine-tuning process. Compared to the available model categories, EntGPT-P follows the zero-shot setting and EntGPT-I uses instruction tuning (on AIDA only) and can be considered as a supervised model.
From fine-tuning only on the AIDA dataset, we observed that EntGPT-I achieves SOTA micro-F1 performance on AQUAINT, ACE2004, CWEB, and WIKI, with AIDA and MSNBC as an exception. Moreover, EntGPT-I achieved an overall mean improvement of +2.1% micro-$F_{1}$ score across all six datasets. At the same time, it also achieved comparable performance against the CoherentED model, which was fine-tuned on data from Wiki and AIDA and outperforms other models. However, it is important to note the significant difference in performance on Wiki dataset; CoherentED achieves a micro-$F_{1}$ score of 0.90, while EntGPT-I (GPT3.5) lags behind significantly, scoring only 0.79. This can be attributed to the fact that CoherentED was fine-tuned on both the Wiki and AIDA datasets, whereas EntGPT-I is only trained on AIDA.
Second, we observe that EntGPT-P outperforms the other models across six different ED benchmarks on a zero-shot setting. EntGPT-P reports an average micro-$F_{1}$ improvement of +7.0% over the second-best model i.e. GPT3.5. Third, our experiments indicate that both EntGPT-P and EntGPT-I can help the large language models generalize better and more accurate information. Compared to the vanilla models i.e. GPT3.5 and Llama2, EntGPT-P and EntGPT-I demonstrate substantial performance improvements across the datasets. As highlighted in Table 2, vanilla Llama2 reports the average micro-$F_{1}$ score of 0.40, EntGPT-P and EntGPT-I with a Llama2 backbone significantly improves the performance of the underlying model with the score of 0.67, and 0.64 respectively. As for the GPT3.5, both the proposed methods boost the performance of the vanilla model, and achieve the SOTA performance within their categories.
Table 3. Test micro-$F_{1}$ scores on five ED benchmarks. We compare EntGPT-P and EntGPT-I with other open-sourced ED methods. Highest micro-$F_{1}$ scores are highlighted in bold, second highest micro-$F_{1}$ scores are underlined. The experiments are re-computed from source code to compare entity names for evaluation.
Model KORE OKE15 OKE16 REU RSS avg
REL 0.618 0.705 0.749 0.662 0.680 0.682
End2End 0.569 0.767 0.783 0.677 0.720 0.703
GENRE 0.542 0.640 0.708 0.697 0.708 0.659
ReFinED 0.567 0.781 0.794 0.680 0.708 0.706
GPT3.5 0.707 0.696 0.687 0.688 0.767 0.709
EntGPT-P (GPT3.5) 0.716 0.767 0.770 0.785 0.808 0.769
EntGPT-I (GPT3.5) 0.753 0.825 0.819 0.808 0.825 0.806
We further evaluated our approach on five entity benchmarks from General Entity Annotator Benchmarking Framework (GERBIL) (Usbeck et al., 2015). Because previous work does not evaluate on these datasets, we could only consider open-source entity disambiguation models for comparison. We found that both EntGPT-I and EntGPT-P achieved improvement over the other SOTA models as shown in Table 3. EntGPT-I and EntGPT-P produce an average micro-$F_{1}$ improvement of +10.0%, and +6.3% respectively over the currently highest performing approach (ReFinED). We take special note of the performance gain of the zero-shot EntGPT-P model, which highlights the effectiveness of the multi-step prompting design.
Table 4. Accuracy on six QA datasets. Most accurate zero-shot scores are highlighted in bold, second most accurate zero-shot scores are underlined. Supervised learning methods which are fine-tuned on each dataset individually. † indicates that the results are obtained directly from the source paper.
Model ARC-C OBQA ANLI CSQA PIQA SiQA
supervised ROBERTa (Liu et al., 2019) 0.430 0.649 0.827 0.805 0.794 0.759
QAGNN (Yasunaga et al., 2021) 0.444 0.678 0.830 0.734 0.796 0.757
GreaseLM (Zhang et al., 2022a) 0.447 0.669 0.833 0.742 0.796 0.833
DRAGON (Yasunaga et al., 2022) 0.486 0.720 0.840 0.760 0.811 0.768
zero-shot LLAMA-65B (Touvron et al., 2023a) 0.560 0.602 - - 0.828 0.523
FLAN-137B (Wei et al., 2021) 0.631 0.784 - - 0.805 -
Llama-2-70B (Touvron et al., 2023b) 0.574 0.602 - - 0.828 0.507
Llama-2-70B-chat 0.613 0.648 0.720 0.590 0.670 0.689
GPT3.5 0.822 0.772 0.710 0.654 0.801 0.695
EntGPT-I (GPT3.5) 0.840 0.804 0.754 0.700 0.830 0.706
4.4.2. Question Answering (RQ1 and RQ2)
We also extended our study to the question answering task to assess the generalizability of the instruction-tuned model and its effectiveness in minimizing hallucination. Using six commonsense reasoning benchmarks, we compared the EntGPT models against existing supervised and zero-shot approaches. The accuracy of these models is reported in Table 4.
We observed that EntGPT-I not only attains SOTA performance on six QA datasets in a zero-shot setting but also achieves comparable, and sometimes superior, performance compared to supervised methods. To be specific, EntGPT-I outperforms the supervised DRAGON model on ARC-C, OBQA, and PiQA datasets, with 35.4%, 8.4%, and 1.9% absolute accuracy gain respectively. In addition, we compared the EntGPT-I model against the foundation GPT3.5 and Llama2 models. We found that the EntGPT-I model outperforms the foundation models. These observations further underscore the effectiveness of leveraging large language models with entity-based knowledge to mitigate hallucination.
4.5. Ablation Study (RQ3)
Table 5. Ablation study on 8 datasets with GPT3.5 backbone. The highest scores are highlighted in bold.
Ablation KORE OKE15 OKE16 REU RSS ACE04 MSN AQU avg
EntGPT-P w/o Aug. (step 2) 0.707 0.696 0.687 0.688 0.767 0.853 0.821 0.753 0.747
EntGPT-P w/o BLINK (step 1) 0.722 0.769 0.748 0.676 0.794 0.890 0.878 0.865 0.793
EntGPT-P 0.716 0.767 0.770 0.785 0.808 0.918 0.867 0.791 0.803
EntGPT-I w/o candidates 0.688 0.717 0.714 0.739 0.765 0.861 0.832 0.732 0.756
EntGPT-I 0.753 0.825 0.819 0.808 0.825 0.937 0.922 0.906 0.849
When investigating the performance of multi-step systems like EntGPT, it is important to study its various parts. Here we show the results of an ablation experiment on eight of the ten entity disambiguation benchmarks – the two very large datasets WIKI and CWEB could not be considered in the ablation study due to budget constraints. This ablation study compared the effectiveness of the entity candidates generation (Step-1) on both EntGPT-P and EntGPT-I, as well as the importance of the augmentation by auxiliary content (Step-2) on EntGPT-P. Note that without Step-1, EntGPT-I should no longer be considered an instruction-tuned model but rather as a fine-tuned model.
Results of the ablation experiment are shown in Table 5 (top). We found that removal of the BLINK entity candidate generation model from the candidates set of EntGPT-P (i.e., use of only the Prior model) did not substantively impact performance. We only found a mean micro-$F_{1}$ improvement of +1.0% across eight benchmarks with the inclusion of BLINK candidates. We also observed that EntGPT-P performed better without BLINK candidates on KORE50, OKE15, MSN and AQU datasets. This may be because BLINK could introduce noise into the entity candidate set, which could cause confusion and hinder performance. Moreover, we found that removing the auxiliary content (i.e. Step-2) severely decreases the performance of EntGPT-P. This evidence shows that auxiliary content enhances the connections between the mention and the target entity.
To further elucidate the importance of grounding entities to reducing hallucination, we fine-tuned GPT3.5 on the AIDA dataset without any entity candidates. Table5 (bottom) shows that removal of these entities produced a sharp decline in the mean micro-$F_{1}$ indicating that this entity-grounding is directly responsible for the performance increase.
Table 6. Comparison of the ground truth label and the label predicted by the EntGPT-P model. In many error cases the EntGPT-P model produces labels that, in our opinion, are more accurate than the labeled ground truth. In other cases where the ground truth was indeed correct, the errant prediction was close.
Dataset Degree of Error Ground Truth (GT) Prediction Source of Error
ACE04
high
Ministry of Defense (Iran)
Ministry of Defense (Japan)
Step 3
ACE04
low
President of Egypt
President
Step 3
ACE04
low
Gaza City
Gaza Strip
Step 2
ACE04
low
Volvo
Volvo Cars
Step 3
KORE
low
First Ladies of Argentina
First Lady
Step 2
KORE
high
Justin Bieber
Justin I
Step 2
KORE
high
Lady Gaga
Gwen Stefani
Step 3
KORE
high
Paul Allen
NULL
Step 2
AQU
high
Cancer
Lung Cancer
Step 3
AQU
high
Tissue (biology)
Facial tissue
Step 3
CWEB
high
Head
Head (company)
Step 2
CWEB
none
Hillsborough County, Florida
Hillsborough, North Carolina
GT is incorrect
CWEB
low
Lake Wylie
Lake Wylie, South Carolina
Step 3
CWEB
none
Australia Cricket Team
Australia
GT is incorrect
MSN
none
New York City
New York
GT is incorrect
MSN
none
University of Alabama
Alabama Crimson Tide football
GT is incorrect
MSN
high
World Trade Center
Collapse of the World Trade Center
Step 3
OKE15
none
Fellow
Research Fellow
GT is incorrect
OKE15
low
Cambridge
University of Cambridge
Step 2
OKE15
none
Principal (academia)
Head teacher
GT is incorrect
OKE15
low
Faculty (academic staff)
Professor
Step 2
OKE15
none
Officer
Officer (armed forces)
GT is incorrect
OKE16
none
Director (business)
Executive director
GT is incorrect
OKE16
none
Germany
Nazi Germany
GT is incorrect
OKE16
none
Czechs
Czech Republic
GT is incorrect
OKE16
none
Sorbonne
University of Paris
GT is incorrect
REU
none
Georgia Power
Georgia (U.S. state)
GT is incorrect
REU
low
Lloyds Bank of Canada
Lloyds Bank
Step 3
RSS
none
Steve Jobs
Apple Inc.
GT is incorrect
RSS
low
Pro Bowl
Super Bowl
Step 3
RSS
none
Eric Kearney
Cincinnati
GT is incorrect
RSS
high
Cleveland Browns
Cleveland
Step 3
5. A Case Study on Entity Disambiguation
In this section, we show the results of a case study of EntGPT’s prediction errors to determine the nature of its mistakes. In general, error-cases were of two forms: (1) incorrect ground truth, and (2) predictions were wrong.
In the first case, we identified cases where the predicted label was, in our opinion, more-correct than the ground truth itself. These cases are indicated at “GT is incorrect” in Table 6. We estimate about one-third of the error can be accounted by errors in the ground truth.
In other cases, EntGPT-P did make mistakes. However, we found that in many of these cases these mistakes were still reasonable predictions as these predictions partially matched the ground truth. For example, in the case where the ground truth is the President of Egypt (line 2 in Table 6), the prediction of “President” was reasonable but too broad. We further analyzed the source of these errors and found that most of the mistakes were caused by the multiple choice selection step (Step 3). The augmentation step (Step 2) occasionally failed to provide useful auxiliary content.
Figure 2. Examples from the entity disambiguation task. We compared GPT3.5, EntGPT-P and EntGPT-I on the OKE15 dataset. Mentions are highlighted within the brackets.
Three example cases from the OKE15 dataset are compared across three models (GPT3.5, EntGPT-P, and EntGPT-I) on disambiguation task are illustrated in Fig. 2. Given the task of linking the mention “Farnborough”, with context provided, EntGPT-I correctly identified “Farnborough, Hampshire”. Yet there were also instances where EntGPT-I failed to correctly link the mention to the desired entity.
6. Case Study on Question Answering (RQ4)
To better understand the results entity linking provides to LLMs, we also conducted a brief error analysis on two of the QA datasets: ARC-C and OBQA. ARC-C is comprised of science exam questions collected from various sources; OBQA is a collection of multiple-choice elementary-level science questions. As reported in Table 4, we performed our experiments with three different models: Llama2-chat, GPT3.5, and EntGPT-I with a GPT3.5 foundation model. We highlight some of the interesting cases as observed in Fig. 3 to assess how different models behave under varying conditions.
Although we do not delve into a comprehensive examination of errors and their various types, we use these specific instances to provide some analysis and context. We found that all of the models fail on mathematical reasoning: e.g., “An astronaut drops a 1.0kg $\ldots$”, which requires an understanding of kinetic energy and the use of some logical deduction. We therefore conclude that enhancing the entity-based knowledge does not necessarily help the LLMs in mathematical reasoning.
Figure 3. Examples from two QA tasks: ARC-C (left), and OBQA (right), showing the performance of different language models on a particular question. Here, GPT refers to the foundational GPT3.5 model, and Llama2 refers to the Llama2-70B-chat.
Moreover, we performed an analysis on cases where the questions required both commonsense and logical reasoning. We found that entity-based knowledge provided the models with a better understanding of the context and therefore helps to mitigate hallucination, with some exceptions. As highlighted in Fig 3, the examples that require both commonsense knowledge and logical deduction such as the predators-prey relationship and the shark example do benefit from grounding the entities. However, we did sometimes encounter instances when entity linking did not improve performance. For example, in the soil case, the entity-linking lead the model astray, finding “roots” rather than the more-direct correct answer “microscope”. Although these errors are unfortunate, they still leave the door open for additional improvement in future work.
7. Discussion
In this work, we described the Entity GPT model and two variants: prompt-based (EntGPT-P) and instruction-tuned (EntGPT-I). By linking natural language mentions of concrete entities generated by large language models to their corresponding entities in some knowledge base, we showed that the EntGPT approach reduces hallucination. Furthermore, we also conclude from our extensive experiments that enhancing entity-based knowledge not only helps in entity disambiguation task but also provides better performance in question answering tasks.
This study also motivates us to explore two different follow-up experiments as future work. We first wish to extend the EntGPT framework to perform entity linking (not just entity disambiguation). Compared to entity disambiguation, entity linking not only requires a model to select correct entities but also to extract mentions. Although previous studies (Cao et al., 2021; Xiao et al., 2023a) have explored the application of generative models to entity linking, they do not consistently demonstrate advantages over supervised methods. Second, we wish to dive deeper into how and why entity disambiguation improves question answering performance. These initial observations point to entity-correlation as a crucial factor leading to the better performance, so it may be wise to incorporate entity correlation more deeply into the question answering paradigm.