### v3

# EntGPT: Entity Linking with Generative Large Language Models

Abstract
Entity Linking (EL) in natural language processing seeks to match text entities to their corresponding entries in a dictionary or knowledge base. Traditional approaches rely on contextual models, which can be complex, hard to train, and poorly transferable across domains. Generative large language models (LLMs) like GPT offer a promising alternative but often underperform with naive prompts. In this study, we introduce EntGPT, employing advanced prompt engineering to enhance EL tasks. Our three-step hard-prompting method (EntGPT-P) significantly boosts the micro-$F_{1}$ score by up to 36% over vanilla prompts, achieving competitive performance across 10 datasets without supervised fine-tuning (SFT). Additionally, our instruction tuning method (EntGPT-I) improves micro-$F_{1}$ scores by 2.1% on average in supervised EL tasks and outperforms several baseline models in six Question Answering (QA) tasks. Our methods are compatible with both open-source and proprietary LLMs. All data and code are available on GitHub at https://github.com/yifding/In_Context_EL.
keywords:
Information Extraction, Generative Large Language Models, Entity Linking
journal: Knowledge-Based Systems
[…]
[label1]organization=University of Notre Dame, city=Notre Dame, state=Indiana, country=USA
[…]
1 Introduction
Figure 1: Pipeline of EntGPT frameworks: Given input document with the annotated mention, both the approaches first conducts (1) entity candidate generation step to obtain relevant entities. EntGPT-Prompting (EntGPT-P) then performs (2) an augmentation step to obtain an auxiliary content of the annotated mention followed by (3) a multi-choice selection prompt to decide the corresponding entity of the annotated mention. EntGPT-Instruction tuning (EntGPT-I), on the other hand, goes through the (2) construction of the instruction dataset, and is (3) instruction-tuned on the curated data.
The advent of Large Language Models (LLMs) has generated significant excitement and concern in the fields of natural language processing (NLP) and artificial intelligence (AI). While their capabilities in text generation and unstructured reasoning are impressive Raffel-JMLR'20-T5; Radford-OpenAI_Blog'18-GPT; Radford-OpenAI_Blog'19-GPT2; Brown-NIPS'20-GPT3; Touvron-arXiv'23-Llama2, their ability to produce structured output remains underdeveloped and relatively unexplored Zhu-AAAI'22-JAKET; Dong-arXiv'23-Head-to-Tail. Entity disambiguation (ED), a crucial task in information extraction (IE), involves linking text fragments representing real-world entities to entries in structured knowledge bases like encyclopedias or dictionaries. These links can enhance various downstream processes by leveraging existing data and relationships within the knowledge base. Integrating deep neural network models, such as Transformers Vaswani-NIPS'17-attention with the symbolic reasoning capabilities of knowledge bases has been termed the “Third Wave of AI”Darpa-Darpa'18-Next_AI. This convergence is essential for advancing problem-solving and reasoning in AI systemsBrooks-AI'81-Symbolic; Hitzler-Book'22-Neuro. For this vision to materialize, it is critical that LLMs are effectively linked to the symbolic entities they reference.
Recent research efforts Qin-arXiv'23-Chatgpt_NLP_Task; Gonzalez-arXiv'23-ChatGPT_Historical_Entities; wei2023zero; li2023prompting have utilized LLMs, specifically GPT-3.5, for IE tasks. However, these studies primarily focus on named entity recognition (NER) and relation extraction (RE) tasks having only a few dozen entity types, which is far less than the 6 million entities present in general Wikipedia entries for the entity disambiguation (ED) task. The current state-of-the-art in ED relies heavily on entity features, predicting entity labels (such as Wikipedia pages or DBpedia entries) from models trained on Wikipedia links Ganea-EMNLP'17-deep_ed, Web hyperlinks Wu-EMNLP'20-BLINK, or info-boxes Ayoola-NAACL'22-ReFinED; Bhargav-NAACL'22-Dbpedia_Entity_Type, and considering the internal structure among entities Hu-KBS'20-Graph_Entity. These methods have been successful in closed-world scenarios where data is clean and complete Hoffart-EMNLP'11-CONLL03_entity, and in open-world scenarios where data may be incomplete but well-described Logeswaran-ACL'19-zero; Wu-EMNLP'20-BLINK. Another approach to the ED task is through text generation. One of the earliest generative models for ED is GENRE DeCao-ICLR'21-GENDRE, which proposed a sequence-to-sequence framework that generates an entity-label sequence from a mention-sequence conditioned with special indicators.
In this work, we use LLMs for the ED task. Most errors in ED arise from rare entities, but auxiliary knowledge from LLMs can help address these challenging cases. Furthermore, most ED models require a ranking step to select the best entity from several candidates within a given context. Generative LLMs have shown exceptional ability in classification tasks xu2023large; fei2022lasuie; josifoski2022genie and can naturally fit into the entity ranking step with their extensive background knowledge.
As illustrated in Fig. 1, we explore two distinct approaches for tackling the ED task with LLMs. The first approach is a multi-step prompting framework called EntGPT-Prompting (EntGPT-P). This method first generates a set of entity candidates for mentions in a document. Then, it uses the LLM to generate auxiliary content, aiding the final selection of the correct entity from the candidate set. Results across various tasks and datasets show that EntGPT-P achieves performance comparable to supervised models without requiring training or fine-tuning on human-annotated data.
The second approach is called EntGPT-Instruction Tuning (EntGPT-I). This method is trained further on instructions (i.e., prompt-response pairs) and demonstrates more controllable and predictable behavior zhang2023instruction. EntGPT-I significantly outperformed EntGPT-P in micro-$F_{1}$ scores across ED benchmarks. Additionally, we evaluated EntGPT-I on six closed-world question-answering tasks. Experiments showed that EntGPT-I achieved state-of-the-art (SOTA) performance in the zero-shot setting, significantly improving accuracy scores over the most popular commercial LLMs.
In summary, our major contributions are as follows:
  • 1.
    We introduce EntGPT-P, a model that follows a three-step hard prompting method to guide generative language models. Our findings demonstrate that EntGPT-P achieves comparable, and sometimes better performance when compared to supervised counterparts across different benchmarks on the ED task.
  • 2.
    We present EntGPT-I, a model instruction-tuned on generative LLMs with instruction-response pairs for the ED task, while augmenting the instructions with enhanced knowledge. We achieve SOTA performance on the ED task.
  • 3.
    We show that these knowledge-enhanced generative models help to generate factually correct answers compared to commercial LLMs through six different Question Answering tasks. EntGPT-I achieves SOTA performance on all datasets in zero-shot setting and comparable performance to supervised methods.
  • 4.
    We also provide a detailed performance comparison of both EntGPT approaches on open-source (e.g., Llama2) and closed-source (e.g., GPT-3.5) LLMs. We find a consistent boost in micro-$F_{1}$/accuracy performance when compared to their vanilla counterparts.
Throughout the remainder of the paper, unless explicitly stated otherwise, we use the following abbreviations: GPT3.5 for gpt-3.5-turbo, Llama for Llama-65B, Llama2 for Llama2-70B, and Llama2-chat for its chat variant Llama2-70b-chat.
2 Related Work
2.1 Entity Disambiguation
Existing research in ED is primarily divided into two categories: feature extraction and task refinement. Key work in feature extraction includes the interaction between mentions and context Ganea-EMNLP'17-deep_ed; Kolitsas-CONLL'18-end2end, the consistency between entities and entity types Raiman-AAAI'18-deeptype; Tedeschi-EMNLP'21-NER4EL; Raiman-AAAI'22-deeptype2, the relationship between entities and knowledge base entries Ayoola-NAACL'22-kb; Ayoola-NAACL'22-ReFinED, and correlation between entities Phan-TKDE'19-Pair_link; Xiao-EMNLP'23-CoherentED. Traditional ED models framed this task as entity classification Ganea-EMNLP'17-deep_ed, while more recent work approached it as machine reading comprehension Barba-ACL'22-Extend, dense retrieval Wu-EMNLP'20-BLINK, question answering Zhang-ICLR'22-Entqa, sequence-to-sequence generation DeCao-ICLR'21-GENDRE, and prompt LLMs Xiao-EMNLP'23-Instruction.
2.2 Prompt-based Methods for Information Extraction
One of the important developments in the field of prompt-based methods for information extraction was the release of GPT-3, which has been shown to outperform state-of-the-art methods for language modeling Brown-NIPS'20-GPT3. The benefit of prompt-based methods is the ability to query a language model to obtain additional knowledge and understanding about a given context, thereby significantly improving performance in zero-shot and few-shot settings Li-arXiv'23-GPT-NER; Dong-arXiv'23-Head-to-Tail. There are mainly three streams of work in this paradigm. The first is in-context learning where a few demonstrative examples are provided to an LLM in order to obtain some desired output. The second is prompt-tuning, where a set of learnable parameters are inserted into the model Lester-EMNLP'21-power; Ding-ACL'22-Prompt_Entity_Typing or each layer Li-ACL'21-prefix; Wang-EMNLP'23-aprompt for fine-tuning. The third line of work is instruction-tuning, where (instruction, response) pairs are prepared used to fine-tune the LLM Longpre-arXiv'23-flan.
In terms of task formulation, existing work on prompt-based methods for information extraction can be primarily categorized into three paradigms: type-oriented, span-oriented, and generation. Type-oriented tasks Ding-ACL'22-Prompt_Entity_Typing mainly locate the given class for a certain mention within the original documents. Span-oriented methods Cui-ACL'21-template_NER_BART enumerate the possible spans and assign corresponding class labels. Finally, generation aims to transform information extraction into a sequence generation schema Li-arXiv'23-GPT-NER.
3 EntGPT
Here, we formally define the ED task and describe the general pipeline employed in both the prompt-based and instruction-tuning approaches to ground generative LLMs with entities from a knowledge base.
3.1 Problem Definition
Given a document represented as a sequence of tokens $\mathcal{D}$ and a set of subsequences $M=\{m_{1},\cdots,m_{n}\}$ (i.e., mentions) contained in $\mathcal{D}$. The goal of the ED task is to establish a mapping for each mention $m_{i}\in M$ to its corresponding entity $e$ in the set $\mathcal{E}$ representing entities in some knowledge base (such as Wikipedia or DBpedia).
3.2 EntGPT-P: Entity Disambiguation with Multi-Step Prompting
We view the task of grounding (or linking) generative LLMs to knowledge bases as a natural use case of ED where text sequences need to be matched with the relevant knowledge base entries to permit the knowledge base’s well-curated structure to be leveraged for downstream tasks.
The EntGPT-Prompting (EntGPT-P) model introduced in this section formulates the ED task as having three steps: 1) Generate and filter entity candidates from knowledge base $\mathcal{E}$; 2) Augment each mention by extracting relevant information with a prompt for a LLM; 3) Combine candidates from Step 1 with context from Step 2, forming a multi-choice question prompt for a LLM.
3.2.1 Step 1: Entity Candidate Generation
Given a mention $m$ in document $D$, the first step is to find a subset of entity candidates $\mathcal{E}_{c}$ corresponding to $m$ within the knowledge base $\mathcal{E}$. We combine two strategies to generate entity candidates. First, we use statistical information of hyperlinks in the knowledge base to surface the most likely entity candidates $\mathcal{E}_{p}$ from $\mathcal{E}$ that exhibit syntactic similarity to the mentions within the document Ganea-EMNLP'17-deep_ed. However, this probabilistic approach, we call the Prior probability-approach or simply Prior, suffers from low recall because it only considers syntactic similarity. In low-recall cases, we additionally employ a dense retrieval model as a backup to augment the entity candidate generation. Specifically, we use the Bidirectional Encoder Representations from Transformers (BERT)-based entity linking (BLINK) model, which primarily uses cleaned Wikipedia hyperlinks to as its retrieval model Wu-EMNLP'20-BLINK, to generate extra entity candidates $\mathcal{E}_{r}$. Therefore, the final entity candidates set from step 1 is $\mathcal{E}_{c}=\mathcal{E}_{p}\cup\mathcal{E}_{r}$ from which we select the top-10 candidates.
3.2.2 Step 2: Augmentation by Prompting
Because many candidates are syntactically similar, distinguishing them is challenging without contextual information. So we augment mentions with relevant information by asking the LLM “What does {mention} represent?” For example, from Fig. 1 (at left), we generate auxiliary content $\mathcal{A}$ of George Bush by asking the LLM “What does George Bush represent?”. We found that generating context information in this way has the following two advantages: 1) The content generated by the LLM considers contextual information from $\mathcal{D}$ and is supplemented with the world knowledge encoded within the LLM. 2) The auxiliary content produced by the system is more specifically targeted towards our task and therefore minimizes the impact of superfluous information.
3.2.3 Step 3: Multiple-choice Selection by Prompting
Armed with the entity candidates set $\mathcal{E}_{c}$ (from Step 1) and the auxiliary content $\mathcal{A}$ (from Step 2), the goal of Step 3 is to select the most appropriate $e\in\mathcal{E}_{c}$. As illustrated in Fig. 1 (at left), we employ an instruct-based prompt to direct the LLM to make a selection from $\mathcal{E}_{c}$ using the auxiliary content $\mathcal{A}$. To distinguish among these entity candidates effectively, we include the first sentence extracted from the knowledge base entry of each entity candidate as a descriptive reference. Also note that each of the entity candidates come from the subset $\mathcal{E}_{c}$ obtained in Step 1, rather than the complete knowledge base $E$. Therefore, this multi-choice setting can not address the situation where the correct entity is not included in $\mathcal{E}_{c}$. To accommodate this situation, we include the option “None” as a catch-all choice for handling such cases.
3.3 EntGPT-I: Entity Disambiguation with Instruction Tuning
In an initial study, described below, we found that EntGPT-P substantially reduces hallucination and helped to deliver more accurate results compared to non-augmented foundation models (see metrics in Tables 4 and  5). Nonetheless, EntGPT-P still under-performed supervised methods indicating room for improvement. To that end, we constructed an additional instruction-tuning model on top of generative LLMs through similar prompts as that of EntGPT-P. In this section, we discuss the processes involved in creating the EntGPT-I model including the construction of an instruction dataset and the instruction-tuning procedure.
3.3.1 Instruction Dataset Construction
An instruction dataset usually consists of three different elements: an instruction, an optional input, and an anticipated output ji2023survey. Usually in the form of natural language, an instruction provides the description of the task to the model, whereas any optional input can provide additional context. In the present work, we curate one of the largest manually annotated entity linking and entity disambiguation datasets, i.e., Accurate online Disambiguation of Entities in Text and Tables (AIDA) hoffart2011robust into the required format (instruction, input, output) needed to fine-tune the an LLM model. An example of such a training instance is shown in Fig. 1 (at right).
3.3.2 Instruction Tuning
We fine-tune two generative models (GPT3.5 and Llama2-chat) in a fully supervised manner on the curated instruction dataset. We used the OpenAI API platform kublik2023gpt to fine-tune GPT3.5, and we used Replicate API [^1] https://replicate.com/docs/get-started/python to fine-tune Llama2-chat. Fine-tuning was performed for only one epoch and we kept all other hyperparameters at their default settings. Although hyper-parameter tuning could potentially enhance performance, hyper-parameter search is outside the scope of this study. After fine-tuning, EntGPT-I is then evaluated on both the ED task and the question answering task.
4 Experiments
To test the two models described in the present work, we used ten public entity disambiguation benchmark datasets and six question answering datasets widely used in the literature. The entity disambiguation task was to link mentions in the dataset with the correct entity in a knowledge base; the question answering task was to determine the correct answer from multiple choices. Using these experiments our goal is to address the following research questions (RQs):
  • 1.
    RQ1: Does multi-step prompting (EntGPT-P) and instruction tuning (EntGPT-I) produce more accurate entity disambiguation output?
  • 2.
    RQ2: Does fine-tuning generative LLMs with auxiliary information gleaned from the linked knowledge base improve accuracy?
  • 3.
    RQ3: How do the various steps involved in the model contribute to the performance of EntGPT?
  • 4.
    RQ4: Do correlations among entities in the knowledge base help to mitigate hallucinations in large language models?
To answer these questions directly, we first introduce the datasets that we will use and several SOTA models.
4.1 Datasets
Following general ED settings introduced in prior work Guo-Semantic_Web'16-Random_Walk; Kolitsas-CONLL'18-end2end; Ayoola-NAACL'22-ReFinED; Barba-ACL'22-Extend, we evaluated our models on six ED datsets: AIDA Milne-CIKM'08-Learning_to_Link, MSNBC (MSN) Cucerzan-EMNLP'07-MSNBC, Aquaint text corpus (AQUAINT/AQU) Milne-CIKM'08-Learning_to_Link, a multilingual training corpus ACE2004 (ACE04) Ratinov-ACL'11-Local_Global_Disambiguation, Walking Named Entity Disambiguation (WNED-WIKI) Guo-Semantic_Web'16-Random_Walk, and WNED-CWEB (CWEB) Gabrilovich-Dataset'13-Clueweb. We also consider five additional datasets from GERBIL Usbeck-WWW'15'-GERBIL: Keyphrase Overlap Relatedness for Entity Disambiguation (KORE50/KORE) Hoffart-CIKM'12-KORE, Open knowledge extraction challenge (OKE-2015/OKE15) Nuzzolese-SWEC'15-OKE15, OKE-2016 (OKE16) Nuzzolese-SWEC'16-OKE16, N3-Reuters-128 (REU) Roder-LREC'14-n3, and N3-RSS-500 (RSS) Roder-LREC'14-n3. We provide the statistics of these benchmarks, including the number of documents, and mentions in Table 1. Each of the experiments used Wikipedia as the background knowledge base. Following the example of Guo and Barbosa (Guo-Semantic_Web'16-Random_Walk), we manually removed spurious mentions that do not appear in the knowledge base as well as repeated documents and empty documents without any mentions.
We found that EntGPT-I (with GPT3.5 as the foundation model) outperformed other approaches on the average micro-F1 score (Tables 4 and  5), so we additionally evaluated EntGPT-I (GPT3.5) on six different commonsense reasoning (question answering) benchmarks: AI2 Reasoning Challenge-Challenge Set (ARC-C) clark2018think, OpenbbookQA (OBQA) mihaylov2018can, Abductive Natural Language Inference (aNLI) bhagavatula2019abductive, CommonsenseQA (CSQA) talmor2018commonsenseqa, Physical Interaction QA (PIQA) bisk2020piqa, and Social Interaction QA (SIQA) sap2019socialiqa.
4.2 Comparison Models
For ED, we categorize the comparison models into three different groups according to their fine-tuning data: Wikipedia and AIDA (Wiki+AIDA), AIDA-only (AIDA), and no fine-tuning (zero-shot). We compared the performance of our models against the following models:
  • 1.
    End2End Kolitsas-CONLL'18-end2end: The End-to-End Neural Entity Linking (End2End) is a global ED and entity linking system based on Deep-ED Ganea-EMNLP'17-deep_ed, but with a more robust recurrent neural network (RNN) architecture that considers mention-context and mention-mention interactions.
  • 2.
    REL Hulst-SIGIR'20-REL: The Radboud Entity Linker (REL) is an ED and entity linking package which combines the NER method FLAIR Akbik-NAACL'19-FLAIR and ED method mulrel-nel Le-ACL'18-mulrel-nel.
  • 3.
    GENRE DeCao-ICLR'21-GENDRE: The Generative Entity REtrieval model (GENRE) is a generation-based model that re-casts the ED task as an entity name generation process.
  • 4.
    ner4el Tedeschi-EMNLP'21-NER4EL: The Named Entity Recognition for Entity Linking (NER4EL) model uses named entity recognition to perform the entity linking task, which is similar to the entity disambiguation task.
  • 5.
    ReFinED Ayoola-NAACL'22-ReFinED: The Representation and Fine-grained typing for Entity Disambiguation (ReFinED) model is an efficient end-to-end zero-shot-capable EL model is a recent system that considers Wikipedia entries as extra features to the ED and entity linking tasks.
  • 6.
    ExtEnd barba2022extend: The Extractive Entity Disambiguation (ExtEnd) model is a transformer-based framework where the disambiguation task is recast as a machine reading comprehension task.
  • 7.
    CoherentED Xiao-EMNLP'23-CoherentED: The Coherent Entity Disambiguation (CoherentED) model aims to enhance the coherence of entity predictions and achieves state-of-the-art performance on most datasets.
In the question answering task we categorize the baselines into two different groups: supervised and zero-shot. In the supervised setting, we follow common experimental methods to report the performance for the following models.
  • 1.
    ROBERTa liu2019roberta: The Robustly Optimized BERT Pretraining Approach (ROBERTa) is an optimized version of the Bidirectional Encoder Representations from Transformers (BERT) devlin2018bert model, which uses Masked Language Modeling (MLM) as its pre-training objective. The pre-trained model is fine-tuned on each question answering benchmark in a fully supervised manner.
  • 2.
    QAGNN & GreaseLM yasunaga2021qa; zhang2022greaselm: The Question Answering Graph Neural Network (QAGNN) and the Graph REASoning Enhanced Language Model (GreaseLM) both fine-tune an LLM using a knowledge graph (KG) on downstream tasks. These architectures are broadly referred to as KG-augmented question answering models.
  • 3.
    DRAGON yasunaga2022deep: Deep Bidirectional Language-Knowledge Graph Pretraining (DRAGON) uses a knowledge base to enhance the pre-training stage and also performs fine-tuning on downstream tasks in a supervised manner.
In the zero-shot setting, we use the following LLMs as foundation models: Large Language Model Meta AI (Llama, Llama2, & Llama2-chattouvron2023llama; Touvron-arXiv'23-Llama2, the Fine-tuned LAnguage Net (FLAN-137B) wei2021finetuned, and the Generative Pre-trained Transformer (GPT3.5) Ouyang-NIPS'22-InstructGPT.
Table 1: Statistics for 11 experimental datasets.
# of Docs # of Mentions
AIDA 231 4485
MSN 20 656
AQU 50 727
ACE04 35 257
CWEB 320 11154
WIKI 319 6793
KORE 50 144
OKE15 101 536
OKE16 173 288
REU 113 650
RSS 357 524
Table 2: Examples from the entity disambiguation task. We compared GPT3.5, EntGPT-P and EntGPT-I on the OKE15 dataset. Mentions are underlined with blue color.
{tabularx}
@ p.52 X @ Example Model Prediction
Blackett spent ten years working at the Cavendish Laboratory as an experimental physicist with Ernest Rutherford and in 1923 became a fellow of King’s college, Cambridge.
GPT3.5: IEEE
EntGPT-P: None
EntGPT-I: Fellow
The man spent time at the Royal Aircraft Establishment in Farnborough, where he made a major contribution to the design of the Mark XIV bomb sight.
GPT3.5: Royal Aircraft Establishment
EntGPT-P: Royal Aircraft Establishment
EntGPT-I: Farnborough, Hampshire
Here Lutuli was confirmed in the Methodist Church and became a lay preacher.
GPT3.5: Methodist Church of Great Britain
EntGPT-P: United Methodist Church
EntGPT-I: Methodism
Table 3: Examples from two QA tasks: ARC-C (left), and OBQA (right), showing the performance of different language models on a particular question. Here, GPT3.5 refers to the foundational GPT3.5 model, and LLAMA2 refers to the Llama2-70B-chat.
Example Model Prediction
ARC-C An astronaut drops a 1.0 kg object and a 5.0 kg object on the Moon. Both objects fall a total distance of 2.0 m vertically. Which of the following best describes the objects after they have fallen a distance of 1.0 m. (A) They have each lost kinetic energy.
(B) They have each gained the same amount of potential energy.
(C) They have each lost the same amount of potential energy.
(D) They have each gained one-half of their maximum kinetic energy
GPT3.5: B EntGPT-I: C LLAMA2: A
In the equatorial Pacific Ocean, upwelling moves cold, deep sea water to the surface. This water is rich in nutrients and dissolved carbon dioxide gas. During an El Niño event, upwelling declines. This causes surface waters to rise in temperature. Which of the following is most likely to occur in the equatorial Pacific Ocean during an El Niño year? (A) decrease in release of carbon dioxide to the atmosphere
(B) increase in photosynthesis by algae in surface waters
(C) increase in dissolved oxygen in surface waters
(D) decrease in precipitation over ocean water
GPT3.5: B EntGPT-I: A LLAMA2: B
OBQA Predators eat:
(A) lions (B) humans (C) bunnies (D) grass
GPT3.5: B EntGPT-I: C LLAMA2: A
The summer solstice in the northern hemisphere is four months before:
(A) May (B) July (C) April (D) October
GPT3.5: B EntGPT-I: B LLAMA2: C
They studied the soil by using:
(A) plants (B) a telescope (C) roots (D) a microscope
GPT3.5: D EntGPT-I: C LLAMA2: D
If you were attacked by a shark and had to punch it sharply where it pulls in air from, you’d use your hand to make contact with: (A) its snout (B) its gills (C) its nose (D) its belly GPT3.5: A EntGPT-I: B LLAMA2: A
4.3 Evaluation Metrics
To maintain a fair comparison across datasets, we use the in-KB micro-F1 score as our evaluation metric for ED following the example of Guo and Barbosa Guo-Semantic_Web'16-Random_Walk. Specifically, being in-KB requires that ground truth mentions correspond to existing knowledge base entries. Empty or out-of-KB entities are removed in the evaluation process. The mean Micro-F1 score is taken per-mention. Although a model may predict non-entities, each mention will have some corresponding ground truth entity. For question answering, we use accuracy as the evaluation metric.
4.4 Results and Findings
In this section, we discuss the results and findings from our study on both the entity disambiguation and question answering benchmarks.
4.4.1 Entity Disambiguation (RQ1 and RQ2)
Table 4: Test micro-F1 scores on standard six benchmarks. With each setting, best scores are highlighted in bold. † indicates that the results are obtained directly from the source paper.
Model AIDA MSN AQU ACE04 CWEB WIKI avg
Wiki+AIDA REL Hulst-SIGIR'20-REL 0.928 0.935 0.873 0.897 0.776 0.780 0.864
End2End Kolitsas-CONLL'18-end2end 0.891 0.933 0.894 0.892 0.760 0.740 0.851
GENRE DeCao-ICLR'21-GENDRE 0.933 0.943 0.899 0.901 0.773 0.874 0.887
ReFinED Ayoola-NAACL'22-ReFinED 0.939 0.941 0.918 0.908 0.794 0.874 0.894
ExtEnD Barba-ACL'22-Extend 0.926 0.947 0.916 0.918 0.777 0.888 0.895
CoherentED Xiao-EMNLP'23-CoherentED 0.894 0.963 0.946 0.934 0.811 0.906 0.909
AIDA GENRE DeCao-ICLR'21-GENDRE 0.886 0.881 0.771 0.823 0.719 0.717 0.795
ner4el Tedeschi-EMNLP'21-NER4EL 0.925 0.892 0.695 0.913 0.685 0.640 0.792
ExtEnd Barba-ACL'22-Extend 0.900 0.945 0.879 0.889 0.766 0.767 0.858
EntGPT-I (Llama2) 0.691 0.744 0.583 0.681 0.556 0.604 0.643
EntGPT-I (GPT3.5) 0.920 0.922 0.906 0.937 0.795 0.791 0.879
zero-shot Llama2 0.432 0.401 0.441 0.381 0.347 0.451 0.408
EntGPT-P (Llama2) 0.708 0.741 0.635 0.746 0.584 0.632 0.674
GPT3.5 0.746 0.821 0.753 0.853 0.638 0.684 0.749
EntGPT-P (GPT3.5) 0.821 0.867 0.791 0.918 0.709 0.808 0.819
A performance comparison over six entity disambiguation benchmarks is presented in Table 4. We categorized these experiments into three groups based on the data used during the fine-tuning process. Compared to the available model categories, EntGPT-P follows the zero-shot setting and EntGPT-I uses instruction tuning (on AIDA only) and can be considered as a supervised model.
From fine-tuning only on the AIDA dataset, we observed that EntGPT-I achieves SOTA micro-F1 performance on AQUAINT, ACE2004, CWEB, and WIKI, with AIDA and MSNBC as an exception. Moreover, EntGPT-I achieved an overall mean improvement of +2.1% micro-$F_{1}$ score across all six datasets. At the same time, it also achieved comparable performance against the CoherentED model, which was fine-tuned on data from Wiki and AIDA and outperforms other models. However, it is important to note the significant difference in performance on Wiki dataset; CoherentED achieves a micro-$F_{1}$ score of 0.90, while EntGPT-I (GPT3.5) lags behind significantly, scoring only 0.79. This can be attributed to the fact that CoherentED was fine-tuned on both the Wiki and AIDA datasets, whereas EntGPT-I is only trained on AIDA.
Second, we observe that EntGPT-P outperforms the other models across six different ED benchmarks on a zero-shot setting. EntGPT-P reports an average micro-$F_{1}$ improvement of +7.0% over the second-best model i.e. GPT3.5. Third, our experiments indicate that both EntGPT-P and EntGPT-I can help the large language models generalize better and more accurate information. Compared to the vanilla models i.e. GPT3.5 and Llama2, EntGPT-P and EntGPT-I demonstrate substantial performance improvements across the datasets. As highlighted in Table 4, vanilla Llama2 reports the average micro-$F_{1}$ score of 0.40, EntGPT-P and EntGPT-I with a Llama2 backbone significantly improves the performance of the underlying model with the score of 0.67, and 0.64 respectively. As for the GPT3.5, both the proposed methods boost the performance of the vanilla model, and achieve the SOTA performance within their categories.
Table 5: Test micro-$F_{1}$ scores on five ED benchmarks. We compare EntGPT-P and EntGPT-I with other open-sourced ED methods. Highest micro-$F_{1}$ scores are highlighted in bold, second highest micro-$F_{1}$ scores are underlined. The experiments are re-computed from source code to compare entity names for evaluation.
Model KORE OKE15 OKE16 REU RSS avg
REL 0.618 0.705 0.749 0.662 0.680 0.682
End2End 0.569 0.767 0.783 0.677 0.720 0.703
GENRE 0.542 0.640 0.708 0.697 0.708 0.659
ReFinED 0.567 0.781 0.794 0.680 0.708 0.706
GPT3.5 0.707 0.696 0.687 0.688 0.767 0.709
EntGPT-P (GPT3.5) 0.716 0.767 0.770 0.785 0.808 0.769
EntGPT-I (GPT3.5) 0.753 0.825 0.819 0.808 0.825 0.806
We further evaluated our approach on five entity benchmarks from General Entity Annotator Benchmarking Framework (GERBIL) Usbeck-WWW'15'-GERBIL. Because previous work does not evaluate on these datasets, we could only consider open-source entity disambiguation models for comparison. We found that both EntGPT-I and EntGPT-P achieved improvement over the other SOTA models as shown in Table 5. EntGPT-I and EntGPT-P produce an average micro-$F_{1}$ improvement of +10.0%, and +6.3% respectively over the currently highest performing approach (ReFinED). We take special note of the performance gain of the zero-shot EntGPT-P model, which highlights the effectiveness of the multi-step prompting design.
Table 6: Accuracy on six QA datasets. Most accurate zero-shot scores are highlighted in bold, second most accurate zero-shot scores are underlined. Supervised learning methods which are fine-tuned on each dataset individually. † indicates that the results are obtained directly from the source paper.
Model ARC-C OBQA ANLI CSQA PIQA SiQA
supervised ROBERTa liu2019roberta 0.430 0.649 0.827 0.805 0.794 0.759
QAGNN yasunaga2021qa 0.444 0.678 0.830 0.734 0.796 0.757
GreaseLM zhang2022greaselm 0.447 0.669 0.833 0.742 0.796 0.833
DRAGON yasunaga2022deep 0.486 0.720 0.840 0.760 0.811 0.768
zero-shot LLAMA-65B touvron2023llama 0.560 0.602 - - 0.828 0.523
FLAN-137B wei2021finetuned 0.631 0.784 - - 0.805 -
Llama-2-70B Touvron-arXiv'23-Llama2 0.574 0.602 - - 0.828 0.507
Llama-2-70B-chat 0.613 0.648 0.720 0.590 0.670 0.689
GPT3.5 0.822 0.772 0.710 0.654 0.801 0.695
EntGPT-I (GPT3.5) 0.840 0.804 0.754 0.700 0.830 0.706
4.4.2 Question Answering (RQ1 and RQ2)
We also extended our study to the question answering task to assess the generalizability of the instruction-tuned model and its effectiveness in minimizing hallucination. Using six commonsense reasoning benchmarks, we compared the EntGPT models against existing supervised and zero-shot approaches. The accuracy of these models is reported in Table 6.
We observed that EntGPT-I not only attains SOTA performance on six QA datasets in a zero-shot setting but also achieves comparable, and sometimes superior, performance compared to supervised methods. To be specific, EntGPT-I outperforms the supervised DRAGON model on ARC-C, OBQA, and PiQA datasets, with 35.4%, 8.4%, and 1.9% absolute accuracy gain respectively. In addition, we compared the EntGPT-I model against the foundation GPT3.5 and Llama2 models. We found that the EntGPT-I model outperforms the foundation models. These observations further underscore the effectiveness of leveraging large language models with entity-based knowledge to mitigate hallucination.
4.5 Ablation Study (RQ3)
Table 7: Ablation study on 8 datasets with GPT3.5 backbone. The highest scores are highlighted in bold.
Ablation KORE OKE15 OKE16 REU RSS ACE04 MSN AQU avg
EntGPT-P w/o Aug. (step 2) 0.707 0.696 0.687 0.688 0.767 0.853 0.821 0.753 0.747
EntGPT-P w/o BLINK (step 1) 0.722 0.769 0.748 0.676 0.794 0.890 0.878 0.865 0.793
EntGPT-P 0.716 0.767 0.770 0.785 0.808 0.918 0.867 0.791 0.803
EntGPT-I w/o candidates 0.688 0.717 0.714 0.739 0.765 0.861 0.832 0.732 0.756
EntGPT-I 0.753 0.825 0.819 0.808 0.825 0.937 0.922 0.906 0.849
When investigating the performance of multi-step systems like EntGPT, it is important to study its various parts. Here we show the results of an ablation experiment on eight of the ten entity disambiguation benchmarks – the two very large datasets WIKI and CWEB could not be considered in the ablation study due to budget constraints. This ablation study compared the effectiveness of the entity candidates generation (Step-1) on both EntGPT-P and EntGPT-I, as well as the importance of the augmentation by auxiliary content (Step-2) on EntGPT-P. Note that without Step-1, EntGPT-I should no longer be considered an instruction-tuned model but rather as a fine-tuned model.
Results of the ablation experiment are shown in Table 7 (top). We found that removal of the BLINK entity candidate generation model from the candidates set of EntGPT-P (i.e., use of only the Prior model) did not substantively impact performance. We only found a mean micro-$F_{1}$ improvement of +1.0% across eight benchmarks with the inclusion of BLINK candidates. We also observed that EntGPT-P performed better without BLINK candidates on KORE50, OKE15, MSN and AQU datasets. This may be because BLINK could introduce noise into the entity candidate set, which could cause confusion and hinder performance. Moreover, we found that removing the auxiliary content (i.e. Step-2) severely decreases the performance of EntGPT-P. This evidence shows that auxiliary content enhances the connections between the mention and the target entity.
To further elucidate the importance of grounding entities to reducing hallucination, we fine-tuned GPT3.5 on the AIDA dataset without any entity candidates. Table7 (bottom) shows that removal of these entities produced a sharp decline in the mean micro-$F_{1}$ indicating that this entity-grounding is directly responsible for the performance increase.
Table 8: Comparison of the ground truth label and the label predicted by the EntGPT-P model. In many error cases the EntGPT-P model produces labels that, in our opinion, are more accurate than the labeled ground truth. In other cases where the ground truth was indeed correct, the errant prediction was close.
Dataset Degree of Error Ground Truth (GT) Prediction Source of Error
ACE04 high Ministry of Defense (Iran) Ministry of Defense (Japan) Step 3
ACE04 low President of Egypt President Step 3
ACE04 low Gaza City Gaza Strip Step 2
ACE04 low Volvo Volvo Cars Step 3
KORE low First Ladies of Argentina First Lady Step 2
KORE high Justin Bieber Justin I Step 2
KORE high Lady Gaga Gwen Stefani Step 3
KORE high Paul Allen NULL Step 2
AQU high Cancer Lung Cancer Step 3
AQU high Tissue (biology) Facial tissue Step 3
CWEB high Head Head (company) Step 2
CWEB none Hillsborough County, Florida Hillsborough, North Carolina GT is incorrect
CWEB low Lake Wylie Lake Wylie, South Carolina Step 3
CWEB none Australia Cricket Team Australia GT is incorrect
MSN none New York City New York GT is incorrect
MSN none University of Alabama Alabama Crimson Tide football GT is incorrect
MSN high World Trade Center Collapse of the World Trade Center Step 3
OKE15 none Fellow Research Fellow GT is incorrect
OKE15 low Cambridge University of Cambridge Step 2
OKE15 none Principal (academia) Head teacher GT is incorrect
OKE15 low Faculty (academic staff) Professor Step 2
OKE15 none Officer Officer (armed forces) GT is incorrect
OKE16 none Director (business) Executive director GT is incorrect
OKE16 none Germany Nazi Germany GT is incorrect
OKE16 none Czechs Czech Republic GT is incorrect
OKE16 none Sorbonne University of Paris GT is incorrect
REU none Georgia Power Georgia (U.S. state) GT is incorrect
REU low Lloyds Bank of Canada Lloyds Bank Step 3
RSS none Steve Jobs Apple Inc. GT is incorrect
RSS low Pro Bowl Super Bowl Step 3
RSS none Eric Kearney Cincinnati GT is incorrect
RSS high Cleveland Browns Cleveland Step 3
5 A Case Study on Entity Disambiguation
In this section, we show the results of a case study of EntGPT’s prediction errors to determine the nature of its mistakes. In general, error-cases were of two forms: (1) incorrect ground truth, and (2) predictions were wrong.
In the first case, we identified cases where the predicted label was, in our opinion, more-correct than the ground truth itself. These cases are indicated at “GT is incorrect” in Table 8. We estimate about one-third of the error can be accounted by errors in the ground truth.
In other cases, EntGPT-P did make mistakes. However, we found that in many of these cases these mistakes were still reasonable predictions as these predictions partially matched the ground truth. For example, in the case where the ground truth is the President of Egypt (line 2 in Table 8), the prediction of “President” was reasonable but too broad. We further analyzed the source of these errors and found that most of the mistakes were caused by the multiple choice selection step (Step 3). The augmentation step (Step 2) occasionally failed to provide useful auxiliary content.
Three example cases from the OKE15 dataset are compared across three models (GPT3.5, EntGPT-P, and EntGPT-I) on disambiguation task are illustrated in Fig. LABEL:ed_case. Given the task of linking the mention “Farnborough”, with context provided, EntGPT-I correctly identified “Farnborough, Hampshire”. Yet there were also instances where EntGPT-I failed to correctly link the mention to the desired entity.
6 Case Study on Question Answering (RQ4)
To better understand the results entity linking provides to LLMs, we also conducted a brief error analysis on two of the QA datasets: ARC-C and OBQA. ARC-C is comprised of science exam questions collected from various sources; OBQA is a collection of multiple-choice elementary-level science questions. As reported in Table 6, we performed our experiments with three different models: Llama2-chat, GPT3.5, and EntGPT-I with a GPT3.5 foundation model. We highlight some of the interesting cases as observed in Fig. LABEL:fig:obqa to assess how different models behave under varying conditions.
Although we do not delve into a comprehensive examination of errors and their various types, we use these specific instances to provide some analysis and context. We found that all of the models fail on mathematical reasoning: e.g., “An astronaut drops a 1.0kg $\ldots$”, which requires an understanding of kinetic energy and the use of some logical deduction. We therefore conclude that enhancing the entity-based knowledge does not necessarily help the LLMs in mathematical reasoning.
Moreover, we performed an analysis on cases where the questions required both commonsense and logical reasoning. We found that entity-based knowledge provided the models with a better understanding of the context and therefore helps to mitigate hallucination, with some exceptions. As highlighted in Fig LABEL:fig:obqa, the examples that require both commonsense knowledge and logical deduction such as the predators-prey relationship and the shark example do benefit from grounding the entities. However, we did sometimes encounter instances when entity linking did not improve performance. For example, in the soil case, the entity-linking lead the model astray, finding “roots” rather than the more-direct correct answer “microscope”. Although these errors are unfortunate, they still leave the door open for additional improvement in future work.
7 Discussion
In this work, we described the Entity GPT model and two variants: prompt-based (EntGPT-P) and instruction-tuned (EntGPT-I). By linking natural language mentions of concrete entities generated by large language models to their corresponding entities in some knowledge base, we showed that the EntGPT approach exhibit exceptional ability in entity disambiguation. Furthermore, we also conclude from our extensive experiments that enhancing entity-based knowledge not only helps in entity disambiguation task but also provides better performance in question answering tasks.
This study also motivates us to explore two different follow-up experiments as future work. We first wish to extend the EntGPT framework to perform entity linking (not just entity disambiguation). Compared to entity disambiguation, entity linking not only requires a model to select correct entities but also to extract mentions. Although previous studies DeCao-ICLR'21-GENDRE; Xiao-EMNLP'23-Instruction have explored the application of generative models to entity linking, they do not consistently demonstrate advantages over supervised methods. Second, we wish to dive deeper into how and why entity disambiguation improves question answering performance. These initial observations point to entity-correlation as a crucial factor leading to the better performance, so it may be wise to incorporate entity correlation more deeply into the question answering paradigm.